A small probe is made of stainless steel and has a length of 60 mm when it stored at a temperature of 18°C.

What is the change in length of the probe when it is being used in an operation and it has warmed up to 36°C? (to 2 s.f and in mm)

[αsteel = 16.0×10−6 K−1]


1
Expert's answer
2021-05-18T11:10:52-0400

The change in length of the probe is


"\Delta L=L_0\alpha\Delta T=(60)(16.0·10^{−6})(36-18)=0.017\text{ mm}."